# First principle prediction of structural distortions in the cuprates and their impact on the electronic structure

### v2

@@ III.3.3 Many-body treatment: dynamic fluctuations
To go beyond band theory, we need a theoretical framework that allows us to include the basic physics of dynamic electronic fluctuations in a computationally efficient manner. As noted above, the fact that the 25% hole-doped BSCCO is paramagnetic (PM) Sachdev (2003); Yoshizaki et al. (1988); Kaiser et al. (2012) argues that each local spin moment fluctuates dynamically (i.e., in time) Zaanen et al. (1996); Lyons et al. (1988); Lyons and Fleury (1988); Brom and Zaanen (2003); Zhu et al. (2023). Prior research has also suggested a close relation between spin fluctuations and high-temperature superconductivity Moriya and Ueda (2000); Zhu et al. (2023). In this section, we perform a many-body study of the system and arrive at two findings: (a) a many-body treatment removes the energetic drive towards a static, inhomogeneous electronic distribution and instead produces a spatially almost uniform static (i.e., time-averaged) spin/charge distribution where the residual spatial inhomogeneity is only introduced by structural distortions; and (b) while many-body effects narrow the electronic bandwidths when compared to the NM DFT calculations, they have little effect on the Fermi surface including the shadow bands.
There are a number of different many-body solid-state methods when one wishes to go beyond mean-field theory. Considering the sizes of the simulation cells we are dealing with even for the simple 2D single-band Hubbard model, we will use computationally efficient slave-boson methods that we have developed recently as both single-site methods Georgescu and Ismail-Beigi (2015, 2017); Georgescu et al. (2021a) and an accurate cluster-based method Jin and Ismail-Beigi (2023). In these approaches, the electron annihilation operator $\hat{c}_{i\sigma}$ is replaced by the product $\hat{f}_{i\sigma}\hat{O}_{i\sigma}$ where we have a chargeless fermion represented by the fermionic annihilation operator (called a spinon) $\hat{f}_{i\sigma}$ and a charged auxiliary or slave boson represented by its lowering operator $\hat{O}_{i\sigma}$. The two subsystems are separated by approximating the ground-state density matrix $\hat{\rho}$ of the combined spinon+boson system by the product form $\hat{\rho}_{f}\times\hat{\rho}_{s}$ where each subsystem has its own density matrix ($\hat{\rho}_{f}$ for spinons and $\hat{\rho}_{s}$ for the slave bosons). The ground state is found by minimizing the total energy
@@
where the auxiliary “magnetic” fields $B_{i\sigma}$ permit static symmetry-breaking solutions of the electronic ground state Georgescu and Ismail-Beigi (2017); Jin and Ismail-Beigi (2023): part of the minimization process is to find the $B_{i\sigma}$ values that minimize the total energy. The spinon Hamiltonian $\hat{H}_{f}$ describes a set of non-interacting electrons moving on a lattice whose tight-binding hoppings are renormalized by expectations values from the slave sector. The spectral properties of the interacting spinon+slave system at the Fermi energy are found by computing them at the Fermi energy of the spinon problem (e.g., by diagonalization of $\hat{H}_{f}$) Riegler et al. (2020). Note that this approach can reproduce the Hartree-Fock predictions if total energy minimization results in $\langle\hat{O}_{i\sigma}^{\dagger}\hat{O}_{j\sigma}\rangle_{s}=1$ and $B_{i\sigma}=U\langle\hat{n}_{i-\sigma}\rangle$, but it has a wider set of solutions with $\langle\hat{O}_{i\sigma}^{\dagger}\hat{O}_{j\sigma}\rangle_{s}\neq 1$ with dynamical spin/charge fluctuations that may have lower energy.
@@
If one enlarges the size of the interacting cluster being treated, the driving force for static symmetry breaking will further weaken, and this is what we find. When we enlarge our interacting problem to involve two neighboring sites and use our cluster slave-boson approach Jin and Ismail-Beigi (2023), we find that energy minimization of the ground states leads to a solution with no static symmetry breaking. Here, we take the DFT-calculated 244-atom $x=0.25$ unit cell electronic structure and perform Wannierization to build the Hubbard model. Therefore, each interacting slave cluster is solved in its own unique local supermodulation crystalline environment as dictated by the extract $\epsilon_{i}$ on-site energies and hoppings $t_{ij}$. Figs. 10(a) and (c) show that the total ground state energy increases when either a static AFM or stripe pattern of magnetic moments is imposed: the lowest energy solution shows no static symmetry-breaking ($B=0$). However, as shown in Fig. 10(b), there are strong AFM correlations between neighboring sites as one expects from the physics of super-exchange. In brief, energetically favored spin-spin correlations are fully dynamic at this doping level. We refer to this solution as a paramagnetic state (PM).
@@
In summary, the reason the DFT NM Fermi surfaces do so well at reproducing the experimental Fermi surfaces lies in the fact that the actual material system at $x=0.25$ has a PM ground state without static local moment Sachdev (2003); Yoshizaki et al. (1988); Scalapino (2012): the DFT NM is the best band theory solution mimicking this fact despite being a high-energy solution within the DFT framework. Of course, the electrons do have significant spin-spin correlations that lower the total energy, but in this case, the correlations are purely dynamic and are not reflected in static spatial symmetry breaking. Hence, while the band structure calculations systematically and sensibly predict low-energy states with spatially inhomogeneous static patterns of charge and spin, the actual system dynamically and quantum mechanically fluctuates among these “snapshots” during the experimental photoemission process leading to an effectively NM-looking final Fermi surface (and this is not the same as averaging over spectra from a large number of static snapshots). Hence, we can now explain the success of the recipe of using NM DFT calculations to describe band structures of cuprate phases that do not display long-range spin/charge order in experiments.
IV Outlook
@@
In conclusion, we have provided a microscopic understanding of the AFM insulating phase of the undoped Bi-2212 system. Additionally, we have uncovered competing stripe orders in the hole-overdoped system, which offers a paradigmatic approach (specifically for BSCCO) to describe stripe orders using DFT without performing complex many-body calculations. Spectroscopically, our non-magnetic DFT band structure calculations remarkably reproduce the observed normal state spectral properties. Furthermore, we have elucidated the structural origin of the ARPES ±$\pm(\pi,\pm\pi)$ and ±$\pm(\pi/4,-\pi/4)$ shadow bands in the hole-overdoped system. Our work underscores the importance of considering the crystal degrees of freedom, including structural distortions, modulations, and realistic oxygen dopant positions, together with state-of-the-art exchange-correlation functional, for an accurate description of various material properties. Since structural distortions do affect electron pairing and superconductivity in cuprates Andersen et al. (2007); McElroy et al. (2005); Foyevtsova et al. (2009); Gastiasoro and Andersen (2018); Nunner et al. (2005); Maśka et al. (2007); Khaliullin et al. (2010); Litak (2009); Rømer et al. (2013); Johnston et al. (2009), this type of approach can serve as the starting point of future research regarding the effect of structural distortions on the fascinating lower-energy physics of the cuprates.
@@
We believe that our study establishes a robust theoretical foundation for more “surgical” structural engineering of cuprates, particularly with regard to manipulating broken translational and rotational symmetries. Moreover, as our DFT ground state captures the dominant low-energy properties in the normal state of Bi-2212, the Wannierized Hamiltonians extracted from our DFT calculations can provide appropriate parameters for model Hamiltonians in future studies of low-energy phenomena in cuprates. As an example, we show in our work that the DFT-derived band structure is in excellent agreement with experiments when computed for using non-magnetic phase. Using many-body calculations, we show that the reason for this agreement is due to the fact that the overdoped material is in a paramagnetic phase with strong dynamical spin fluctuations (which are known to affect superconductivity and the pseudogap Zhu et al. (2023); Moriya and Ueda (2000)) but without any static spatial symmetry breaking of the local moments or charge. These observations also provide guidance to future researchers regarding which DFT predictions can be compared directly to experimental observables and which other predictions require further refinement using many-body correlated electron approaches.
V Computational Details
@@
We use the Vienna ab initio simulation package (VASP) with the projector-augmented wave method Kresse and Joubert (1999). A relatively high plane-wave cutoff energy of 500 eV is used. The generalized-gradient-approximation (GGA) with the semilocal Perdew–Burke–Ernzerhof (PBE) functional Anisimov et al. (1991); Perdew et al. (1996) is used in all of our calculations. All calculations are done with collinear spins. The recent finding of non-collinear spin texture in cuprates Gotlieb et al. (2018) is also an interesting topic to study but is beyond the scope of this work. To avoid the known failure of the local-spin-density approximation (LSDA) and the GGA in reproducing the copper magnetic moment in cuprates Mattheiss and Hamann (1988); Hybertsen and Mattheiss (1988); Lin et al. (2006); Foyevtsova et al. (2010); Song et al. (2019); He et al. (2008); Fan and Zeng (2011) due to self-interaction errors (SIE) in the approximate exchange-correlation functionals Perdew and Zunger (1981), we add $U=4$ eV for the Cu $3d$ manifold in all our PBE+$U$ calculations following previous theoretical works Yelpo et al. (2021); Wang et al. (2006); Deng et al. (2019). Changing the $U$ value only shifts the energy of the unoccupied high-energy Cu-derived bands Jin and Ismail-Beigi which does not affect our main findings around the Fermi energy.
@@
More advanced and superior exchange-correlation functionals such as the strongly-constrained-and-appropriately-normed (SCAN) meta-GGA functional Sun et al. (2015) can be employed, and they usually require smaller $U$ value (as they inherently better remove self-interaction errors). We have performed SCAN+U calculations as well (see the supplement Jin and Ismail-Beigi ) using $U$ values for prior works for cuprates Long et al. (2020). For both undoped and doped Bi-2122, we find that the SCAN+U method Sai Gautam and Carter (2018); Long et al. (2020) provides very similar results to the PBE+U results Jin and Ismail-Beigi . Details on optimized lattice structures as well as typical meta-stable structures can also be found in the supplementary materials Jin and Ismail-Beigi . Given the similarity of the PBE+U and SCAN+U results and considering the increased computational cost of using SCAN+U, most of the results in our paper use the PBE+U approach.
@@
To accelerate the structural relaxation of the large 244-atom oxygen-doped unit cell, we use the Spanish Initiative for Electronic Simulations with Thousands of Atoms (SIESTA) package Soler et al. (2002) to approximately relax the structure before doing the final relaxations using VASP. A DZP basis with EnergyShift of 100 meV and SplitNorm of 0.25 is used in all our SIESTA calculations.
@@
We substantiate this picture quantitatively by computing the tight-binding Kohn-Sham Hamiltonian on the maximally localized Wannier basis Marzari and Vanderbilt (1997) extracted from our DFT calculations using Wannier90 Pizzi et al. (2020).
@@
Our slave-boson calculations use the Wannierized tight-binding model from our DFT calculations to construct the Hubbard model. We perform the single-site slave-boson calculations using the Boson Subsidiary-Solver (BoSS) software Georgescu et al. (2021b). The cluster slave-boson calculations follow early work Jin and Ismail-Beigi (2023).
VI Acknowledgements

### v3

@@ III.3.3 Many-body treatment: dynamic fluctuations
To go beyond band theory, we need a theoretical framework that allows us to include the basic physics of dynamic electronic fluctuations in a computationally efficient manner. As noted above, the fact that the 25% hole-doped BSCCO is paramagnetic (PM) [134, 135, 136] argues that each local spin moment fluctuates dynamically (i.e., in time) [141, 130, 133, 131, 132]. Prior research has also suggested a close relation between spin fluctuations and high-temperature superconductivity [142, 132]. In this section, we perform a many-body study of the system and arrive at two findings: (a) a many-body treatment removes the energetic drive towards a static, inhomogeneous electronic distribution and instead produces a spatially almost uniform static (i.e., time-averaged) spin/charge distribution where the residual spatial inhomogeneity is only introduced by structural distortions; and (b) while many-body effects narrow the electronic bandwidths when compared to the NM DFT calculations, they have little effect on the Fermi surface including the shadow bands.
There are a number of different many-body solid-state methods when one wishes to go beyond mean-field theory. Considering the sizes of the simulation cells we are dealing with even for the simple 2D single-band Hubbard model, we will use computationally efficient slave-boson methods that we have developed recently as both single-site methods [143, 144, 145] and an accurate cluster-based method [146]. In these approaches, the electron annihilation operator $\hat{c}_{i\sigma}$ is replaced by the product $\hat{f}_{i\sigma}\hat{O}_{i\sigma}$ where we have a chargeless fermion represented by the fermionic annihilation operator (called a spinon) $\hat{f}_{i\sigma}$ and a charged auxiliary or slave boson represented by its lowering operator $\hat{O}_{i\sigma}$. The two subsystems are separated by approximating the ground-state density matrix $\hat{\rho}$ of the combined spinon+boson system by the product form $\hat{\rho}_{f}\times\hat{\rho}_{s}$ where each subsystem has its own density matrix ($\hat{\rho}_{f}$ for spinons and $\hat{\rho}_{s}$ for the slave bosons). The ground state is found by minimizing the total energy
@@
where the auxiliary “magnetic” fields $B_{i\sigma}$ permit static symmetry-breaking solutions of the electronic ground state [144, 146]: part of the minimization process is to find the $B_{i\sigma}$ values that minimize the total energy. The spinon Hamiltonian $\hat{H}_{f}$ describes a set of non-interacting electrons moving on a lattice whose tight-binding hoppings are renormalized by expectations values from the slave sector. The spectral properties of the interacting spinon+slave system at the Fermi energy are found by computing them at the Fermi energy of the spinon problem (e.g., by diagonalization of $\hat{H}_{f}$) [147]. Note that this approach can reproduce the Hartree-Fock predictions if total energy minimization results in $\langle\hat{O}_{i\sigma}^{\dagger}\hat{O}_{j\sigma}\rangle_{s}=1$ and $B_{i\sigma}=U\langle\hat{n}_{i-\sigma}\rangle$, but it has a wider set of solutions with $\langle\hat{O}_{i\sigma}^{\dagger}\hat{O}_{j\sigma}\rangle_{s}\neq 1$ with dynamical spin/charge fluctuations that may have lower energy.
@@
If one enlarges the size of the interacting cluster being treated, the driving force for static symmetry breaking will further weaken, and this is what we find. When we enlarge our interacting problem to involve two neighboring sites and use our cluster slave-boson approach [146], we find that energy minimization of the ground states leads to a solution with no static symmetry breaking. Here, we take the DFT-calculated 244-atom $x=0.25$ unit cell electronic structure and perform Wannierization to build the Hubbard model. Therefore, each interacting slave cluster is solved in its own unique local supermodulation crystalline environment as dictated by the extract $\epsilon_{i}$ on-site energies and hoppings $t_{ij}$. Figs. 10(a) and (c) show that the total ground state energy increases when either a static AFM or stripe pattern of magnetic moments is imposed: the lowest energy solution shows no static symmetry-breaking ($B=0$). However, as shown in Fig. 10(b), there are strong AFM correlations between neighboring sites as one expects from the physics of super-exchange. In brief, energetically favored spin-spin correlations are fully dynamic at this doping level. We refer to this solution as a paramagnetic state (PM).
@@
In summary, the reason the DFT NM Fermi surfaces do so well at reproducing the experimental Fermi surfaces lies in the fact that the actual material system at $x=0.25$ has a PM ground state without static local moment [134, 135, 107]: the DFT NM is the best band theory solution mimicking this fact despite being a high-energy solution within the DFT framework. Of course, the electrons do have significant spin-spin correlations that lower the total energy, but in this case, the correlations are purely dynamic and are not reflected in static spatial symmetry breaking. Hence, while the band structure calculations systematically and sensibly predict low-energy states with spatially inhomogeneous static patterns of charge and spin, the actual system dynamically and quantum mechanically fluctuates among these “snapshots” during the experimental photoemission process leading to an effectively NM-looking final Fermi surface (and this is not the same as averaging over spectra from a large number of static snapshots). Hence, we can now explain the success of the recipe of using NM DFT calculations to describe band structures of cuprate phases that do not display long-range spin/charge order in experiments.
IV Modulation of correlation strength
Figure 12: (a) Top view of CuO2 layer: blue balls are Cu, and red balls are O. The black ball shows the projected position of the nearest O dopant between the Bi and Sr layers while the black dashed line marks the $a$-axis position of this dopant. (b) The AFM correlation $\langle\hat{N}_{i\uparrow}\hat{N}_{j\downarrow}+\hat{N}_{i\downarrow}\hat{N}_{j\uparrow}\rangle_{s}$ between nearest-neighbor Cu from the cluster slave-boson calculations (a Cu pair is identified by the position of their linking planar O atom); the O positions along the $a$-axis are in fractions of $a$. Any inhomogeneity along the $b$-axis is averaged. Blue circles show the results from the relaxed $x=0.25$ hole-doped structure which has superlattice modulation (SM). Red diamonds show data for the symmetrized Wannier model which is spatially uniform. (c) The local onsite energy difference $\epsilon_{pd}$ for the DFT NM state. Each data point shows the averaged $dp$-splitting between an oxygen atom and its two copper neighbors. Green triangles are for the frozen structure obtained by removing oxygen dopants from the $x=0.25$ structure, and purple stars are for the relaxed (modulated) undoped $x=0$ structure. (d) The local CuO bond length between an O atom and its two Cu neighbors averaged over the two bonds.
Due to the atomically flat and clean BiO-terminated surfaces that form on BSCCO upon cleaving of crystals, surface tunneling microscopy and spectroscopy can provide a wealth of spatially-resolved electronic information. A striking experimental discovery [36, 37] was a significant modulation of the superconducting gap with local structural motifs such as (inferred) dopant positions or supermodulation. This experimental information is of significant theoretical interest because any microscopic model for superconductivity will have a dependence on the local doping level as well as the local structural details: for example, a leading theory for the microscopic pairing mechanism is through AFM spin correlations and fluctuations [148, 149] and local doping levels and inter-site hopping parameters (which depend on local bond lengths and angles) help determine the magnetic properties.
However, it is very difficult to directly interpret the experimental data without theoretical models. To date, a large number of theoretical works have attempted to describe the situation [38, 39, 40, 41, 42, 43, 44, 45] using model Hamiltonians and perturbative expansions, but there is a lack of consensus on the dominant expansions terms, and a number of key parameters must be fit to experiment. Separately, a number of these works assume that the electrostatic potential of the dopant plays the main physical role, an assumption whose validity is hard to assess without independent and parameter-free theoretical results.
In this section, we use our DFT+many-body calculations to shed light on some key questions (the methodology is the same as Section III.3.3 and use the same PM state below). We begin with Fig. 12(a) that illustrates a top view of the CuO2 layer we focus on and marks the position of the (in-plane projection) of the oxygen dopants (which are at the necking region of the superlattice modulation). Fig. 12(b) shows the many-body results for the modulation of AFM correlation along the superlattice modulation direction ($a$-axis) for the relaxed $x=0.25$ system. The correlation strength is larger closer to the dopant oxygen, which is consistent with experimental observations [36, 37]. For reference, a flat red line marks the AFM correlation from solving a symmetrized Hubbard model with all hopping strengths and onsite energies set to their average value over the unit cell.
Next, we investigate the relation between the AFM correlation and the local onsite energy difference $\epsilon_{dp}$ between the Cu-$d_{x^{2}-y^{2}}$ and O-$p_{\sigma}$ Wannier orbitals (which, as explained above, are found by Wannierization of the NM DFT+U state within a three-band model). Fig. 12(c) shows the local $dp$ splitting $\epsilon_{pd}$, and we see a clear modulation and direct connection to the AFM correlation strength. The connection is easy to understand: larger $\epsilon_{dp}$ leads to a lower effective hopping between nearest-neighbor Cu sites and reduces Cu-Cu magnetic couplings (see, e.g. Ref. [38]). We expect $\epsilon_{pd}$ to be directly connected to the more experimentally relevant “charge transfer gap” $\Delta_{pd}$(between the top of the O-$p$ type hole states and Cu-$d$ type electron states in experiment): for a fixed Cu $U$ and near half-filling, one expects $\Delta_{pd}\approx\epsilon_{pd}+U/2$. In this way, we expect the modulation of $\epsilon_{pd}$ should reflect that of the charge-transfer gap: Fig. 12(c) predicts an ±$\approx\pm 0.1$ eV modulation of the charge-transfer gap which is consistent with STM measurements [69]. In addition, since our results show that smaller $\epsilon_{pd}$ leads to stronger AFM correlation, within the spin fluctuation paring model we would expect this leads to stronger superconducting pairing which is consistent with experimental observations [88, 86].
We now exploit the power of first-principle calculations, namely that they can deliver parameter-free results along with full control of the microscopic structure, to clearly separate the role of doping from that of structural distortions. We first take the fully relaxed $x=0.25$ structure which has dopants and supermodulation and simply remove the oxygen dopants without allowing any atoms to move (i.e., frozen structure). We then recalculate the electronic structure and show the resulting $\epsilon_{pd}$ in green in Fig. 12(c): the main physical effect is that $\epsilon_{pd}$ curve shifts up uniformly in energy upon removal of the dopants without much change of the amplitude of spatial modulation. Hence, the electrostatic effect and doping profile from the oxygen dopants are in fact quite uniform (i.e., non-local) and do not in themselves create any strong spatial modulation of electronic parameters. Obviously, this argues against a key assumption of many prior model Hamiltonian works and helps us focus on other relevant sources for the modulation (i.e., structural properties). The uniform upward shift of $\epsilon_{pd}$ with the removal of hole dopants is easy to understand: the Cu now becomes more negatively charged and thus their onsite energies get raised up in energy. For reference, we also relax the new $x=0$ structure which retains the superlattice modulation but with weaker amplitude: its resulting $\epsilon_{pd}$ in Fig. 12(c) is consistently higher than the $x=0.25$ case but is much flatter than the unrelaxed $x=0$ structure.
While there are many potential structural parameters or descriptors one could investigate, we focus on the simplest and most important one for $\epsilon_{pd}$: the Cu-O bond length. Fig. 12(d) shows the local Cu-O bond length in the CuO2 plane for all the four systems described above. In the necking regions of the supermodulation, which is also where the dopants are located, the local Cu-O bonds are longest which translates into small local $\epsilon_{pd}$. The connection between Cu-O bond length and $\epsilon_{pd}$ is easy to understand, e.g., from classical electrostatics: for larger Cu-O separation, the Madelung electrostatic onsite potential energies are lowered at Cu and raised at O which reduces $\epsilon_{pd}$. The inverse relation of the two is easy to see when comparing comparable curves in Figs. 12(c) and (d).
In short, we have shown quite clearly that local structural modulations are the dominant controllers of the local $dp$ energy splitting and thus the local AFM correlation strength. In turn, this structural effect helps us understand the cause of variations in the charge-transfer gap and the superconducting gap observed in experiments. Furthermore, it highlights the point that controlling and engineering the local structure of the superconducting layers is the most straightforward path to engineering the superconductivity of the cuprates.
V Outlook
@@
In conclusion, we have provided a microscopic understanding of the AFM insulating phase of the undoped Bi-2212 system. Additionally, we have uncovered competing stripe orders in the hole-overdoped system, which offers a paradigmatic approach (specifically for BSCCO) to describe stripe orders using DFT without performing complex many-body calculations. Spectroscopically, our non-magnetic DFT band structure calculations remarkably reproduce the observed normal state spectral properties. Furthermore, we have elucidated the structural origin of the ARPES ±$\pm(\pi,\pm\pi)$ and ±$\pm(\pi/4,-\pi/4)$ shadow bands in the hole-overdoped system. Finally, we show how the dopant oxygens and structural modulations vary the charge-transfer gap and spin correlations over real space: in particular, we show that the local structure around the CuO2 plane is of paramount importance in controlling the local AFM correlation strength and charge-transfer gap, while the effect of the electrostatic doping effect itself is quite minimal. Our work underscores the importance of considering the crystal degrees of freedom, including structural distortions, modulations, and realistic oxygen dopant positions, together with state-of-the-art exchange-correlation functional, for an accurate description of various material properties. A telegraphic summary is “structure is king.”
@@
We believe that our study establishes a robust theoretical foundation for more “surgical” structural engineering of cuprates, particularly with regard to manipulating broken translational and rotational symmetries. Moreover, as our DFT ground state captures the dominant low-energy properties in the normal state of Bi-2212, the Wannierized Hamiltonians extracted from our DFT calculations can provide appropriate parameters for model Hamiltonians in future studies of low-energy phenomena in cuprates. As an example, we show in our work that the DFT-derived band structure is in excellent agreement with experiments when computed for using non-magnetic phases. Using many-body calculations, we show that the reason for this agreement is due to the fact that the overdoped material is in a paramagnetic phase with strong dynamical spin fluctuations (which are known to affect superconductivity and the pseudogap [132, 142]) but without any static spatial symmetry breaking of the local moments or charge. These observations also provide guidance to future researchers regarding which DFT predictions can be compared directly to experimental observables and which other predictions require further refinement using many-body correlated electron approaches.
VI Computational Details
@@
We use the Vienna ab initio simulation package (VASP) with the projector-augmented wave method [150]. A relatively high plane-wave cutoff energy of 500 eV is used. The generalized-gradient-approximation (GGA) with the semilocal Perdew–Burke–Ernzerhof (PBE) functional [151, 152] is used in all of our calculations. All calculations are done with collinear spins. The recent finding of non-collinear spin texture in cuprates [153] is also an interesting topic to study but is beyond the scope of this work. To avoid the known failure of the local-spin-density approximation (LSDA) and the GGA in reproducing the copper magnetic moment in cuprates [154, 155, 81, 77, 92, 68, 91] due to self-interaction errors (SIE) in the approximate exchange-correlation functionals [156], we add $U=4$ eV for the Cu $3d$ manifold in all our PBE+$U$ calculations following previous theoretical works [73, 157, 158]. In the supplementary material [85], we describe further tests that vary $U$ from 0-4 eV. At $U=0$ eV, we reproduce the non-magnetic metallic ground states from prior works [81, 77, 92, 91]. Once $U>2$ eV, changing the $U$ value only shifts the energy of the unoccupied high-energy Cu-derived bands [85] which does not affect our main findings around the Fermi energy.
@@
More advanced and superior exchange-correlation functionals such as the strongly-constrained-and-appropriately-normed (SCAN) meta-GGA functional [159] can be employed, and they usually require smaller $U$ value (as they inherently better remove self-interaction errors). We have performed SCAN+U calculations as well (see the supplement [85]) using $U$ values for prior works for cuprates [160]. For both undoped and doped Bi-2122, we find that the SCAN+U method [161, 160] provides very similar results to the PBE+U results [85]. Details on optimized lattice structures as well as typical meta-stable structures can also be found in the supplementary materials [85]. Given the similarity of the PBE+U and SCAN+U results and considering the increased computational cost of using SCAN+U, most of the results in our paper use the PBE+U approach.
@@
To accelerate the structural relaxation of the large 244-atom oxygen-doped unit cell, we use the Spanish Initiative for Electronic Simulations with Thousands of Atoms (SIESTA) package [162] to approximately relax the structure before doing the final relaxations using VASP. A DZP basis with EnergyShift of 100 meV and SplitNorm of 0.25 is used in all our SIESTA calculations.
@@
We substantiate this picture quantitatively by computing the tight-binding Kohn-Sham Hamiltonian on the maximally localized Wannier basis [83] extracted from our DFT calculations using Wannier90 [84].
@@
Our slave-boson calculations use the Wannierized tight-binding model from our DFT calculations to construct the Hubbard model. We perform the single-site slave-boson calculations using the Boson Subsidiary-Solver (BoSS) software [163]. The cluster slave-boson calculations follow early work [146].
VII Acknowledgements